# arXiv:2603.01660v1[eess.SP] 02 Mar 2026

### v1

@@ III-A Design of IRS Weights
To achieve performance analogous to a bi-static/multi-static radar, the IRS-static radar must incorporate the effective field-of-view (FOV) of the corresponding distributed configuration into its design. Consider a bi-static radar as illustrated in Fig. 2(a), where the intersection of the transmit and receive beams defines the FOV for target detection. In the equivalent IRS-static radar depicted in Fig. 2(b), the bi-static receiver is replaced by the IRS panel. The IRS weights are then designed to redirect the incoming signal toward the radar, which hosts both the transmitter and the receive array.
Figure 2: Field-of-view for (a) Bi-static radar, (b) IRS-static radar.
To realize the desired FOV, the bi-static radar receiver would be oriented toward the direction $\big(\psi^{\left(\mathrm{az}\right)}_{0},\psi^{\left(\mathrm{el}\right)}_{0}\big)$. For the IRS-static radar, the IRS weights must be configured such that signals received from this angular region are redirected toward the radar receiver, which is located at an angle $\big(\psi^{\left(\mathrm{az}\right)}_{\mathrm{R}},\psi^{\left(\mathrm{el}\right)}_{\mathrm{R}}\big)$ relative to the IRS panel. The IRS weight design primarily aims to compensate for the relative phase across the reflecting elements along the overall propagation path. Defining the diagonal matrices $\mathbf{A}_{0}=\mathrm{diag}\{\mathbf{a}(\psi^{\left(\mathrm{az}\right)}_{0},\psi^{\left(\mathrm{el}\right)}_{0})\}$ and $\mathbf{A}_{\mathrm{R}}=\mathrm{diag}\{\mathbf{a}(\psi^{\left(\mathrm{az}\right)}_{\mathrm{R}},\psi^{\left(\mathrm{el}\right)}_{\mathrm{R}})\}$, the diagonal IRS weight matrix $\mathbf{D}$ should satisfy
@@
For targets in the vicinity of $\big(\psi^{\left(\mathrm{az}\right)}_{0},\psi^{\left(\mathrm{el}\right)}_{0}\big)$, the IRS panel redirects the response back to the radar. The maximum beamforming gain from the IRS is achieved when the target is located exactly in the direction $\big(\psi^{\left(\mathrm{az}\right)}_{0},\psi^{\left(\mathrm{el}\right)}_{0}\big)$. The performance of this IRS weight design is analyzed in the simulation results section.
III-B Recasting the signal model for Target Parameter Estimation
@@
It is evident from (7) that the received signal corresponds to a point in 3-D space defined by the intersection of $\mathbf{a}_{L_{v}L_{h}}\left(\psi_{p}^{\left(\mathrm{az}\right)},\psi_{p}^{\left(\mathrm{el}\right)}\right)$ and $\mathbf{a}_{M_{T}}\left(\theta_{p}^{\left(\mathrm{az}\right)}\right)$, further constrained by the round-trip delay. From (7), we combine the total gain ($\alpha_{p}$) and the transmitter steering vector $(\textbf{a}_{MT})$ with the beamformed transmit signal (s) so as to enable estimation of the angle information with respect to the IRS. The resulting compact representation is
$\overline{y}_{k}=\overline{\textbf{F}}\textbf{D}\textbf{A}\overline{\textbf{x}}+\overline{\textbf{v}}_{k},$ (12)
@@
where $\overline{\textbf{x}}=\sum^{P_{k}}_{p=1}\alpha_{p}\textbf{a}_{MT}(\theta^{az}_{p},\theta^{el}_{p})\textbf{s}$ and the dictionary matrix $\mathbf{A}\in\mathbb{C}^{L_{h}L_{v}\times G_{R}}$ corresponds to the AoA space around the IRS such that
$\mathbf{A}=\big[\mathbf{a}(\psi_{1}),\dots,\mathbf{a}(\psi_{G_{R}})\big].$ (13)
@@
The grid points $(\psi_{G_{R}})$ are expressed as ${\Psi}=\{\psi_{z}:\psi_{z}=\left(\psi_{z}^{(\mathrm{az})},\psi_{z}^{(\mathrm{el})}\right),\psi_{z}^{(\mathrm{az})}\in\left[-\pi/2,\pi/2\right],\psi_{z}^{(\mathrm{el})}\in\left[0,\pi/2\right]\ 1\leq z\leq G_{R}\}$. Based on the target field of view with respect to the IRS, the angular region (and hence the number of grid points) can be reduced. In the following section, Cramer-Rao Bounds are derived for the signal model presented above.
IV Cramer Rao Bounds
We identify the target parameters from the proposed signal model (12) as
$\mathbf{\zeta}=[\psi^{\mathrm{az}},\psi^{\mathrm{el}}]^{T},$ (14)
@@ IV Cramer Rao Bounds
where $\psi^{\mathrm{az}}=[\psi^{\mathrm{az}}_{1},\psi^{\mathrm{az}}_{2},\ldots,\psi^{\mathrm{az}}_{k}]$ and $\psi^{\mathrm{el}}=[\psi^{\mathrm{el}}_{1},\psi^{\mathrm{el}}_{2},\ldots,\psi^{\mathrm{el}}_{k}]$ denote the true azimuth and elevation angles, respectively, for $k$ targets, resulting in $\mathbf{\zeta}\in\mathbb{R}^{2k~\text{x}~1}$. Our objective is to characterize the fundamental error bound, i.e., the minimum achievable variance, for estimating these angle parameters under the proposed $k$-target signal model using the Cramér–Rao Lower Bound (CRLB) framework.
Several approaches exist for deriving Cramér–Rao bounds depending on the assumptions and structure of the signal model [11]. For IRS-assisted radar with a beamformed transmit signal (s) treated as an unknown waveform, a corresponding CRB derivation is provided in [5]. In our setting, we assume the signal waveform to be deterministic and the test statistics to be known. Accordingly, we adopt the CRB formulation for known deterministic waveforms for planar arrays presented in [11] in order to obtain the angle-estimation error bounds associated with the IRS observations.
@@
$\textbf{y}(k)=\overline{\textbf{A}}\overline{\textbf{x}}(k)+\overline{\textbf{v}},$ (15)
where $\overline{\textbf{A}}=\overline{\textbf{F}}\textbf{D}\textbf{A}(\psi^{\mathrm{az}},\psi^{\mathrm{el}})$. This compact representation reduces notational complexity by capturing the signal structure across snapshots within a single matrix form.
@@
$\mathbf{J}=\begin{bmatrix}\mathbf{J}_{\psi^{\mathrm{az}},\psi^{\mathrm{az}}}&\mathbf{J}_{\psi^{\mathrm{az}},\psi^{\mathrm{el}}}\\ \mathbf{J}_{\psi^{\mathrm{el}},\psi^{\mathrm{az}}}&\mathbf{J}_{\psi^{\mathrm{el}},\psi^{\mathrm{el}}}\end{bmatrix}$ (16)
The vector derivative of the steering vector $\overline{\textbf{A}}$ is defined as
$\mathbf{Z}_{i}=\left[\frac{\partial\mathbf{\overline{a}}(\boldsymbol{\psi}_{i})}{\partial\psi_{x_{i}}}\quad\frac{\partial\mathbf{\overline{a}}(\boldsymbol{\psi}_{i})}{\partial\psi_{y_{i}}}\right]\quad i=1,\cdots,k,$ (17)
and
$\mathbf{Z}=\left[\mathbf{Z}_{1}\quad\mathbf{Z}_{2}\quad\cdots\quad\mathbf{Z}_{k}\right],$ (18)
which forms an $N\times 2k$ matrix. Here, we define
$\mathbf{H}\triangleq\begin{bmatrix}\mathbf{Z}^{H}_{\psi_{az}}\mathbf{Z}_{\psi_{az}}&\mathbf{Z}^{H}_{\psi_{az}}\mathbf{Z}_{\psi_{el}}\\ \mathbf{Z}^{H}_{\psi_{el}}\mathbf{Z}_{\psi_{az}}&\mathbf{Z}^{H}_{\psi_{el}}\mathbf{Z}_{\psi_{el}}\end{bmatrix}$ (19)
We use the subscript 2 to denote that $\boldsymbol{\psi}_{i}$ has two components. We define a $2\times 2$ matrix of ones as
$\mathbf{1}_{2\times 2}=\begin{bmatrix}1&1\\ 1&1\end{bmatrix}$. (20)
Therefore, the expression for the FIM under i.i.d. snapshots is
$\textbf{J}=\frac{2K}{\sigma_{w}^{2}}\left[Re\left[\mathbf{H}\odot\left[\mathbf{S}_{\mathbf{f}}^{T}\otimes\mathbf{1}_{2\times 2}\right]\right]\right],$ (21)
where $\sigma_{w}^{2}$, $K$, $S_{f}$, $\otimes$, and $\odot$ denote the noise variance, the number of snapshots, the source signal powers, and the Kronecker and Hadamard products, respectively. The resulting CRB matrix is given by
$C_{CR}=J^{-1}$ (22)
$C_{CR}=\begin{bmatrix}\mathbf{C}_{\psi^{\mathrm{az}},\psi^{\mathrm{az}}}&\mathbf{C}_{\psi^{\mathrm{az}},\psi^{\mathrm{el}}}\\ \mathbf{C}_{\psi^{\mathrm{el}},\psi^{\mathrm{az}}}&\mathbf{C}_{\psi^{\mathrm{el}},\psi^{\mathrm{el}}}\end{bmatrix}$ (23)
Consequently, the variance and Root-Mean-Square Error (RMSE) for azimuth and elevation estimates are obtained as
$var(\psi_{az})=C_{\psi_{az},\psi_{az}},RMSE(\psi_{az})=\sqrt{C_{\psi_{az},\psi_{az}}}$ (24)
$var(\psi_{el})=C_{\psi_{el},\psi_{el}},RMSE(\psi_{el})=\sqrt{C_{\psi_{el},\psi_{el}}}$ (25)

### v2

@@ III-A Design of IRS Weights
To achieve performance analogous to a bi-static/multi-static radar, the IRS-static radar must incorporate the effective field-of-view (FOV) of the corresponding distributed configuration into its design. From the IRS-static radar depicted in Fig. 1(c), we note that the bistatic receiver is replaced by the IRS panel. To realize the desired FOV, the bi-static radar receiver would be oriented toward the direction $\big(\psi^{\left(\mathrm{az}\right)}_{0},\psi^{\left(\mathrm{el}\right)}_{0}\big)$. For the IRS-static radar, the IRS weights must be configured such that signals received from this angular region are redirected toward the radar receiver, which is located at an angle $\big(\psi^{\left(\mathrm{az}\right)}_{\mathrm{R}},\psi^{\left(\mathrm{el}\right)}_{\mathrm{R}}\big)$ relative to the IRS panel. The coherent signal arrival at the receive array serves as the design objective for IRS weights. Defining the diagonal matrices $\mathbf{A}_{0}=\mathrm{diag}\{\mathbf{a}(\psi^{\left(\mathrm{az}\right)}_{0},\psi^{\left(\mathrm{el}\right)}_{0})\}$ and $\mathbf{A}_{\mathrm{R}}=\mathrm{diag}\{\mathbf{a}(\psi^{\left(\mathrm{az}\right)}_{\mathrm{R}},\psi^{\left(\mathrm{el}\right)}_{\mathrm{R}})\}$, the diagonal IRS weight matrix $\mathbf{D}$ should satisfy
@@
For targets in the vicinity of $\big(\psi^{\left(\mathrm{az}\right)}_{0},\psi^{\left(\mathrm{el}\right)}_{0}\big)$, the IRS panel redirects the response back to the radar with some beamforming gain, attaining maximum gain when the target is located exactly in the direction $\big(\psi^{\left(\mathrm{az}\right)}_{0},\psi^{\left(\mathrm{el}\right)}_{0}\big)$. Although this design does not incorporate the desired beamwidth that determines the FOV, we employ this for the target-parameter estimation CRB analysis. The performance of this IRS weight design is analyzed in Section V.
III-B Reconstruction of Signal Model
Stacking the received signal in (5) over $C$ chirps, we obtain the radar data cube $\mathbf{\overline{Y}}\in\mathbb{C}^{M\times N\times C}$, where $N$ denotes the number of fast-time samples. Performing the discrete Fourier transform (DFT) along the slow-time dimension, we to obtain $M_{R}$ Range-Doppler (RD) maps, $\mathbf{\overline{Y}}_{\mathrm{RD}}=\mathcal{F}_{C}\{\mathbf{\overline{Y}}\}$, where $\mathcal{F}_{C}\{.\}$ is a $C$-point DFT operator. Due to the superposition of the $K$ target responses, $\mathbf{\overline{Y}}_{\mathrm{RD}}=\sum\limits_{k=1}^{K}\mathbf{\overline{Y}}^{k}_{\mathrm{RD}}$, where $\mathbf{\overline{Y}}^{k}_{\mathrm{RD}}$ denotes the RD contribution due to the $k$th target alone. Using either constant false-alarm-rate (CFAR) detector (or its variants) [9] or technique developed in [3], one can obtain $K_{P}\leq K$ RD bins corresponding to true detections; with some bins potentially containing multiple targets. Therefore, the total number of targets $K=\sum_{p=1}^{K_{P}}P_{k}$, where $P_{k}$ denotes the number of targets present in the $k$th detected RD bin, $(\hat{R}_{k},\hat{v}_{k})$. Note that $\hat{R}_{k}=\frac{c\tau_{k,l}|_{l=0}}{2}$ corresponds to the round-trip distance via the IRS, while $\hat{v}_{k}$ denotes the relative bi-static velocity of the target with respect to the radar and the IRS.
The observation vector for the $k$th RD bin across the $M_{R}$ receivers is given by
$\mathbf{\overline{y}}_{k}=\mathbf{FD}\sum_{p=1}^{P_{k}}\mathbf{h}_{p}\mathbf{g}_{p}^{H}\mathbf{s}+\mathbf{\bar{v}}_{k}.$ (8)
We denote the LoS channel between the radar and the $p$th target as $\mathbf{g}_{p}=\alpha_{p}^{(g)}\mathbf{a}_{M_{T}}(\theta^{(\mathrm{az})}_{p},\theta^{(\mathrm{el})}_{p})$, where $\mathbf{a}_{M_{T}}(\theta^{(\mathrm{az})}_{p},\theta^{(\mathrm{el})}_{p})$ is the steering vector toward the target as seen from the transmit array. Similarly, we write $\mathbf{h}_{p}=\alpha_{p}^{(h)}\mathbf{a}_{L_{h}L_{v}}\left(\psi^{(\mathrm{az})}_{p},\psi^{(\mathrm{el})}_{p}\right)$, where $\mathbf{a}_{L_{h}L_{v}}\left(\psi^{(\mathrm{az})},\psi^{(\mathrm{el})}\right)=\mathbf{a}_{L_{v}}\left(\psi^{(\mathrm{az})}\right)\otimes\mathbf{a}_{L_{h}}\left(\psi^{(\mathrm{el})}\right)$ is the steering vector toward the target as observed from the IRS panel, and $\otimes$ denotes the Kronecker product. Defining $\alpha_{p}=\alpha_{p}^{(h)}\alpha_{p}^{(g)}$, (8) can be rewritten as
$\mathbf{\overline{y}}_{k}=\mathbf{FD}\sum_{p=1}^{P_{k}}\alpha_{p}\mathbf{a}_{L_{h}L_{v}}\left(\psi_{p}^{\left(\mathrm{az}\right)},\psi^{\left(\mathrm{el}\right)}_{p}\right)\\ \times\mathbf{a}_{M_{T}}\left(\theta^{\left(\mathrm{az}\right)}_{p},\theta^{\left(\mathrm{el}\right)}_{p}\right)^{H}\mathbf{\overline{s}}+\mathbf{\bar{v}}_{k}.$ (9)
With reference to Fig. 1(b) for the detection obtained at ($\hat{R}_{k}$$\hat{v}_{k}$), the parameters $\theta^{\left(\mathrm{az}\right)}_{k},\theta^{\left(\mathrm{el}\right)}_{k},\psi^{\left(\mathrm{az}\right)}_{k},\psi^{\left(\mathrm{el}\right)}_{k},\forall\ k\in[1,K]$ are to be estimated. Importantly, the 3-D target position can be determined using only a subset of these parameters. Let $R_{k,\mathrm{R}}$ and $R_{k,\mathrm{I}}$ denote the distances from the $k$th target to the radar and to the IRS, respectively. Then the estimated range inferred from the path delay satisfies
$\hat{R}_{k}=R_{k,\mathrm{R}}+R_{k,\mathrm{I}}+R_{\mathrm{I},\mathrm{R}},$ (10)
where the distance between IRS and the radar $R_{\mathrm{I},\mathrm{R}}$ is assumed known. It follows that estimating $\psi^{(\mathrm{az})}_{k}$ and $\psi^{(\mathrm{el})}_{k}$, together with $R_{k,\mathrm{I}}$, yields the $k$th target position $\mathbf{r}_{k}^{\prime}=[x_{k}^{\prime},y_{k}^{\prime},z_{k}^{\prime}]^{T}$ with respect to the IRS coordinate system. The target position in the radar coordinate system can then be obtained via
$\mathbf{r}_{k}=\mathbf{r}_{k}^{\prime}+\mathbf{r}_{\mathrm{I}}.$ (11)
Alternatively, the target position can be obtained directly by estimating $\theta^{\left(az\right)}_{k}$ and $\theta^{\left(el\right)}_{k}$ along with $R_{k,\mathrm{R}}$ from $\hat{R}_{k}$.
@@ III-B Reconstruction of Signal Model
It is evident from (9) that the received signal corresponds to a point in 3-D space that is defined by the intersection of $\mathbf{a}_{L_{v}L_{h}}\left(\psi_{p}^{\left(\mathrm{az}\right)},\psi_{p}^{\left(\mathrm{el}\right)}\right)$ and $\mathbf{a}_{M_{T}}\left(\theta_{p}^{\left(\mathrm{az}\right)}\right)$, further constrained by the round-trip delay. (9) can be compactly represented as:
$\overline{\textbf{y}}_{k}=\textbf{F}\textbf{D}\textbf{A}\overline{\textbf{x}}+\overline{\textbf{v}}_{k},$ (12)
@@
where $\overline{\textbf{x}}=\sum^{P_{k}}_{p=1}\alpha_{p}\textbf{a}_{MT}(\theta^{az}_{p},\theta^{el}_{p})\textbf{s}$ and the dictionary matrix $\mathbf{A}\in\mathbb{C}^{L\times G_{R}}$ corresponds to the angle-of-arrival space around the IRS such that
$\mathbf{A}=\big[\mathbf{a}(\psi_{1}),\dots,\mathbf{a}(\psi_{G_{R}})\big],$ (13)
@@
where the angular grid points $\psi_{1},...,\psi_{G_{R}}]$ are expressed as ${\Psi}=\{\psi_{z}:\psi_{z}=\left(\psi_{z}^{(\mathrm{az})},\psi_{z}^{(\mathrm{el})}\right),\psi_{z}^{(\mathrm{az})}\in\left[-\pi/2,\pi/2\right],\psi_{z}^{(\mathrm{el})}\in\left[0,\pi/2\right]\ 1\leq z\leq G_{R}\}$. Based on the target field of view with respect to the IRS, the angular region (and hence the number of grid points, $G_{R}$) can be reduced. In the following section, Cramer-Rao Bounds are derived for the signal model presented above.
IV Cramér-Rao Bounds
The target parameters to be estimated from all detected observation vectors $\overline{\mathbf{y}}_{k}$ are
$\boldsymbol{\zeta}$ $=[\zeta_{1},\zeta_{2},\ldots,\zeta_{K}],$ (14)
@@ IV Cramér-Rao Bounds
where, $\boldsymbol{\zeta}_{i}=[\psi_{i}^{(\mathrm{az})},\,\psi_{i}^{(\mathrm{el})}]^{T}$is the target parameters correponding to $k$-th target and $\mathbf{\zeta}\in\mathbb{R}^{2\times K}$ is the target parameters for all $K$ targets as seen from the IRS. Our objective is to characterize the fundamental error bound, i.e., the minimum achievable error variance, for estimating these angle parameters under the proposed signal model using the Cramér–Rao Bound (CRB) framework.
Several approaches exist for deriving CRB depending on the assumptions and the structure of the signal model [11]. In other IRS-assisted radar configurations, CRB has been derived for signal waveform estimation in [5] and for target parameters estimation in [17]. In this work, we derive CRB for target parameter estimation assuming deterministic transmit signal waveform in (P3) IRS-assisted configuration. We develop CRB for IRS-based $\zeta_{i},i\in[1,K]$, which is equivalent to the study w.r.t. the radar because of the positional equivalence observed in (11).
@@
$\overline{\textbf{y}}_{k}=\overline{\textbf{A}}\overline{\textbf{x}}+\overline{\textbf{v}},$ (15)
where $\overline{\textbf{A}}=\textbf{F}\textbf{D}\textbf{A}$. This compact form of the signal model ensures that only the target parameters as seen from the IRS, i.e., $(\psi^{\mathrm{(az)}},\psi^{\mathrm{(el)}})$, are explicit while the rest are embedded within $\overline{\mathbf{A}}$.
@@
$\small\mathbf{J}=\begin{bmatrix}\mathbf{J}_{\psi^{(\mathrm{az})},\psi^{(\mathrm{az})}}&\mathbf{J}_{\psi^{(\mathrm{az})},\psi^{(\mathrm{el})}}\\ \mathbf{J}_{\psi^{(\mathrm{el})},\psi^{(\mathrm{az})}}&\mathbf{J}_{\psi^{(\mathrm{el})},\psi^{(\mathrm{el})}}\end{bmatrix}$. (16)
To obtain $\mathbf{J}$, we first define the vector derivative for each column of $\overline{\textbf{A}}$ with respect to target parameters as
$\mathbf{Z}_{i}=\left[\frac{\partial\mathbf{\overline{a}}(\boldsymbol{\zeta}_{i})}{\partial\psi_{i}^{(\mathrm{az})}}\quad\frac{\partial\mathbf{\overline{a}}(\boldsymbol{\zeta}_{i})}{\partial\psi_{i}^{(\mathrm{el})}}\right]\quad\in\mathbb{C}^{M_{R}\times 2},i=1,\cdots,K,$ (17)
Stacking these derivatives of all target parameters for $K$ targets, we get
$\mathbf{Z}=\left[\mathbf{Z}_{1}\quad\mathbf{Z}_{2}\quad\cdots\quad\mathbf{Z}_{K}\right]\quad\in\mathbb{C}^{M_{R}\times 2K},$ (18)
Defining $\textbf{H}=\textbf{Z}^{H}\textbf{P}_{\frac{1}{\overline{\textbf{A}}}}\textbf{Z},$, where $\textbf{P}_{\frac{1}{\overline{\textbf{A}}}}=\textbf{I}-\overline{\textbf{A}}(\overline{\textbf{A}}^{H}\overline{\textbf{A}})^{-1}\overline{\textbf{A}}^{H}$ is the orthogonal projection matrix of $\overline{\textbf{A}}$, (16) can be compactly expressed as [11]:
$\textbf{J = }\frac{2N}{\sigma_{w}^{2}}\,\mathrm{Re}\!\left[\mathbf{H}\odot\left(\mathbf{S}_{\mathbf{f}}^{T}\otimes\mathbf{1}_{2\times 2}\right)\right],$ (19)
where $\sigma_{w}^{2}$, $N$, $S_{f}$ and $\mathbf{1}_{2\times 2}$ denote the noise variance, the number of snapshots, the source signal powers and $2\times 2$ matrix of ones, respectively. $\odot$ is the Hadamard product operator. The resulting CRB expression is given by
$\textbf{C}_{CR}=\textbf{J}^{-1}=\begin{bmatrix}\mathbf{C}_{\psi^{\mathrm{(az)}},\psi^{\mathrm{(az)}}}&\mathbf{C}_{\psi^{\mathrm{(az)}},\psi^{\mathrm{(el)}}}\\ \mathbf{C}_{\psi^{\mathrm{(el)}},\psi^{\mathrm{(az)}}}&\mathbf{C}_{\psi^{\mathrm{(el)}},\psi^{\mathrm{(el)}}}\end{bmatrix}$. (20)
The $\mathbf{C}_{CR}$ contains the variance corresponding to the azimuth and elevation of each target. Consequently, the Root-Mean-Square Error (RMSE) for azimuth and elevation estimates are obtained as
$RMSE(\psi^{(\mathrm{az})})=\sqrt{\textbf{C}_{\psi^{(\mathrm{az})},\psi^{(\mathrm{az})}}},$ (21)
$RMSE(\psi^{(\mathrm{el})})=\sqrt{\textbf{C}^{\psi^{(\mathrm{el})},\psi^{(\mathrm{el})}}}.$ (22)
Although the derived CRB resembles the standard form of a linear signal model, the critical contribution of the IRS path to the performance of the IRS-static radar configuration is embedded within the CRB. Besides the dependence of CRB on the snapshots ($N$) and the SNR $(\sigma^{2}_{w},S_{f})$, the matrix $H$ embeds the system setup in relation with target positions as seen from the IRS. Since the considered IRS-static configuration is new, we employ the derived CRB to analyse the sensitivity/robustness of this configuration across various target and system scenarios for the target parameter estimation in the subsequent section.